let vector a=i+j+k and let r be a variable vector such that r.i , r.jand r.k are positive integers. if r.a<=12 then what are the no. of values of r?

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear shefali

 let r =xi +yj +zk


r.i =x ≥1

r.j=y  ≥1

r.k=z  ≥1

and r.a =x+y+z ≤12

  x+y+z ≤12

let u=x-1

   v=y-1

  w=z-1

 then   u+v+w ≤9     where u≥0,v≥0,w≥0

now we have to find the intiger solution of this equation.

let  u+v+w =n   where n varies from 0 to 9

  solution of these equation

 n:0 to 9  ∑n+3-1C3-1

=n:0 to 9  ∑n+2C2

=n:0 to 9  ∑(n+2)(n+1)/2

=n:0 to 9  1/2 ∑(n2+3n+2)

= 1/2 [9(9+1)(18+1)/6   + 3*9(9+1)/2  + 2 *10]

 =1/2[285 +135+20] =220

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin